OpenStudy (anonymous):

You drop a ball from a height of 1.5 meters. Each curved path has 71% of the height of the previous path. a. Write a rule for the sequence using centimeters. The initial height is given by the term n = 1. b. What height will the ball be at the top of the sixth path?

OpenStudy (anonymous):

A(n)=150(71/100)^n-1 Convert (1.5m=150cm) n=6 A(6)=150(71/100)^6-1 A(6)=150(0.71)^5 A(6)=150(0.180) A(6)=27.06 cm